The outer edge of a rotating Frisbee with a diameter of 30cm has a linear speed of 3.2m/s . What is the angular speed of the Frisbee? Express your answer using two significant figures. I got 10.6 rounded to 11 radian/sec but its saying its wrong... am i missing something? I took 3.2/.3m...
Answer #1
angular speed of the Frisbee= V/r = 3.2 /0.15 =21.3333 rad/s